# Diffusion Models Through a Global Lens: Are They Culturally Inclusive?

### v1

@@ Abstract
Text-to-image diffusion models have recently enabled the creation of visually compelling, detailed images from textual prompts. However, their ability to accurately represent various cultural nuances remains an open question. In our work, we introduce CultDiff benchmark, evaluating state-of-the-art diffusion models whether they can generate culturally specific images spanning ten countries. We show that these models often fail to generate cultural artifacts in architecture, clothing, and food, especially for underrepresented country regions, by conducting a fine-grained analysis of different similarity aspects, revealing significant disparities in cultural relevance, description fidelity, and realism compared to real-world reference images. With the collected human evaluations, we develop a neural-based image-image similarity metric, namely, CultDiff-S, to predict human judgment on real and generated images with cultural artifacts. Our work highlights the need for more inclusive generative AI systems and equitable dataset representation over a wide range of cultures.
@@
Zahra Bayramlithanks: Equal contribution   Ayhan Suleymanzade11footnotemark: 1   Na Min An   Huzama Ahmad Eunsu Kim   Junyeong Park   James Thornethanks: Corresponding author   Alice Oh22footnotemark: 2 KAIST
@@ 1 Introduction
Text-to-image (T2I) diffusion models recently have shown a significant advance in generating high-quality images from text prompts(Saharia et al., 2022; Ramesh et al., 2022; Betker et al., 2023; Podell et al., 2023; Blattmann et al., 2023; Esser et al., 2024). Remarkable success has been demonstrated in producing realistic and detailed visual representations, ranging from generating complex scenes Liu et al. (2024b) to performing style transfer using diffusion models Wang et al. (2023); Zhang et al. (2023). While it is crucial to address the challenge of generating culture-aware imagery, current diffusion models have yet to be evaluated on their ability to create a distribution of accurate and fair representations of images Cao et al. (2024b); Liu et al. (2023).
A recent study, Basu et al. (2023) reveals that while current diffusion models can generate realistic and contextually relevant images, they frequently fail to achieve geographical inclusiveness, generating well-represented images worldwide. Furthermore, a critical gap remains in these models’ ability to produce culturally accurate representations, failing to create concepts or artifacts for particular cultures, thus, lacking in culture awareness Kannen et al. (2024). While benchmark datasets have been developed to capture cultural diversity Jha et al. (2023); Myung et al. (2024); Li et al. (2024c), to the best of our knowledge, there is limited prior work exploring the ability of diffusion models to generate culturally specific artifacts, particularly those representing underrepresented cultures Kannen et al. (2024); Nayak et al. (2024).
@@ 2.1 Cultural-Aware Datasets
Beginning with the benchmark to evaluate image quality from simpler text prompts using text-to-image (T2I) diffusion models, such as HPDv2 Wu et al. (2023) using human preferences, more recently advanced efforts include creating comprehensive benchmarks such as T2I-CompBench Huang et al. (2023) and GenAI-Bench Li et al. (2024a). These prior benchmarks assess the performance of T2I models across various aspects, including realism, fidelity, and compositional text-to-image generation. Although a very recent study Jha et al. (2024) has started investigating global representation in T2I models, their emphasis has primarily been on regional social stereotypes.
@@
On the other hand, many widely-used datasets for training T2I synthesis models, such as LAION-400M Schuhmann et al. (2021), tend to exhibit Anglo-centric and Euro-centric biases, as noted by Birhane et al. (2021). These biases skew the representation of cultures in generated images, often favoring Western perspectives. In response to this, several researchers have worked to create datasets that better represent diverse cultures. For example, the MaRVL dataset Liu et al. (2021) was specifically designed to include a broader array of languages and cultural concepts, covering regions such as Indonesia, Swahili-speaking Africa, Tamil-speaking South Asia, Turkey, and Mandarin-speaking China and addressing biases in datasets that predominantly focused on North American and Western European cultures.
@@
In the Large Language Models (LLMs) domains, SeeGULL dataset Jha et al. (2023) broadens stereotype benchmarking to encompass global and regional diversity, and BLEnD benchmark Myung et al. (2024) evaluates the cultural knowledge of LLMs across various languages including low-resource ones. Similarly, the Dollar Street dataset Gaviria Rojas et al. (2022) sought to capture everyday life across a wide variety of socioeconomic backgrounds, presenting a more globally inclusive view. In addition, Liu et al. (2023) introduced the CCUB dataset developed to promote cultural inclusivity by collecting images and captions representing the cultural contexts of different countries. Most recently, CUBE Kannen et al. (2024) is a large-scale dataset of cultural artifacts spanning 8 countries across different geo-cultural regions for evaluating cultural diversity. Our work further contributes to dataset creation by expanding the focus to include low-resource cultures, thereby addressing gaps between overrepresented and underrepresented cultural contexts.
@@ 3.1 CultDiff Dataset Construction
To acquire real images for training and evaluation, we collected images for each artifact across all categories and countries from the internet using the generated prompts. Recognizing that real-world images vary significantly due to factors such as location, camera angle, lighting, and weather conditions, we aimed to minimize potential biases by selecting five images rather than solely one for each artifact. The images were scraped using Bulk Bing Image Downloader (BBID) [^4] https://github.com/ostrolucky/Bulk-Bing-Image-downloader, an open-source Python tool that enables fast downloads and filters out adult content.
@@
3.2 Human Evaluation
A critical part of our experimental design involved human annotation to assess the cultural and contextual relevance of AI-generated images. We first created a survey comprising 150 images for all ten countries (e.g., United States, China, Azerbaijan, Indonesia), where the images for each country were equally divided across three categories: architecture, clothing, and food items. Each survey included a set of four images for each artifact: three real images (collected via web scraping) and one AI-generated image (produced by one of the three diffusion models).
To prevent model bias, each human annotator was distributed with randomly selected images generated from one of three diffusion models for each question. Three human annotators from each country participated in the study, ensuring cultural familiarity and context-specific insights. All 30 annotators were asked to evaluate each image based on the following multiple-choice questions:
Image-image similarity scores
are collected for each category as follows:
Image-text alignment scores
Realism scores
are assessed based on how realistic the AI-generated image appears.
3.3 Model Evaluation
3.3.1 Pair and Dataset Creation
@@ 3.3.2 Similarity Metric
To build a model-based image-image similarity metric, CultDiff-S, we adopted a contrastive learning approach, utilizing Vision Transformers (ViT) (Alexey, 2020) to learn embeddings for both real and AI-generated images, as illustrated in Figure 2. The global attention mechanism in ViT enables effective "local-to-global" correspondence, enhancing self-supervised learning (Caron et al., 2021; Mo et al., 2023). This ViT property is particularly important in our work, where it helps evaluate the cultural and contextual relevance of images by capturing both local features and their broader context. To achieve this, we employed a weighted margin loss during the training process, utilizing the scores (collected from our human annotation process explained in Section 3.2) as weights ($w_{i}$, $i=1,2,...,N$ in the equation below, motivated by previous studies that utilized GPT feedback scores as the weights in building sentence embedding models Cheng et al. (2023); An et al. (2024). These scores are normalized to 0 to 1 and applied to their corresponding real-synthetic (human-annotated) positive and negative image pairs. For real-image pairs that were not annotated by humans, we assigned a default weight of $w=1$. Our margin loss is defined as:
$\mathcal{L}\;=\;\frac{1}{N}\sum_{i=1}^{N}\;w_{i}\,\Bigl{(}y_{i}\,d_{i}^{2}\;+\;(1-y_{i})\,\max\bigl{(}0,m-d_{i}\bigr{)}^{2}\Bigr{)}$
@@ 4.1.2 Image-Description Match
Our analysis focused on the responses to the second question (Q2), which assessed the fidelity of the generated content to the provided description. Figure 3 presents the results for the FLUX model’s outputs across architecture, clothing, and food, respectively (refer to Appendix A.3 for results from other models).
The United States and the United Kingdom consistently achieved the highest average fidelity scores across all three categories. In contrast, Azerbaijan, South Korea, and Ethiopia recorded the lowest average scores of 2.28, 1.55, and 1.54 in the architecture, food, and clothing categories, respectively. Nevertheless, these average scores reveal a noticeable gap between the scores of the United States and those of other countries across all categories. These results suggest a potential bias in the model, favoring countries with a more extensive cultural presence on the internet. Addressing this limitation by incorporating more diverse and balanced datasets could improve the model’s fidelity for underrepresented countries and enhance its overall cultural accuracy Ananthram et al. (2024); Liu et al. (2025).
@@ 4.1.3 Realism
4.2 Similarity Metric Analysis
After training the similarity model, we evaluated its performance using our test dataset pairs. For each evaluation pair, we computed additional similarity metrics, including Fréchet Inception Distance (FID) (Heusel et al., 2017), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), and the Structural Similarity Index Measure (SSIM) (Wang et al., 2004). Additionally, we calculated the correlation between image-image similarity scores—derived by averaging the responses to the first four questions in the survey—and each of these similarity metrics. In our metric model, we compute the cosine similarity of the embeddings produced by the model during evaluation as the final similarity score. The results of these correlations are summarized in Table 2. The proposed metric, CultDiff-S, achieves notably higher correlation values, suggesting the usefulness of using our automatic pre-trained image-image metric, showing potential to replace the previous metrics.
@@ 5.1 Human-Aligned Image-Image Similarity Evaluation Metric
Designing an image-image similarity metric that closely correlates with human rankings is inherently difficult due to the complexity of cultural attributes and subjective interpretations. Culture itself is a broad and multifaceted concept, making it challenging to define in a universally accepted manner Liu et al. (2024a). In fact, many previous studies in related domains that incorporate cultural aspects either provide only a high-level definition or avoid in-depth discussions on its nuances Adilazuarda et al. (2024); Huang and Yang (2023); Zhou et al. (2023); Cao et al. (2024a); Wan et al. (2023). In our study, we treated countries as representatives of distinct cultures. However, individuals, even within the same country, may have diverse interpretations of cultural artifacts, reflecting regional, ethnic, and personal differences in cultural perception, which can lead to subjective similarity scores given by human annotators. Additionally, cultural artifacts often evolve over time, influenced by various internal or external factors, further complicating the task of cultural representations. As a result, an effective similarity metric that correlates more closely with human perception must account for both objective attributes, such as structural and visual features, and subjective factors, including historical context, regional variations, and personal interpretations. We believe that developing such a metric necessitates integrating human feedback into the process, combining both data-driven approaches and human-centered evaluations to bridge the gap, as has been demonstrated in other contexts and domains in prior works Xu et al. (2023); Stiennon et al. (2020); Ouyang et al. (2022), as well as in our research.
@@ A.1 Prompt Details
  • “A panoramic view of {landmark} in {country}, realistic”
  • “An image of {clothes} from {country} clothing, realistic”
  • “An image of {food} from {country} cuisine, realistic”
@@
  • “A panoramic view of the Empire State Building in the United States, realistic”
  • “An image of Hanfu from Chinese clothing, realistic”
  • “An image of plov from Azerbaijani cuisine, realistic”
@@
A.2 Survey Question Details
This section provides additional details regarding the survey questions and instructions used during the study. Annotators received clear guidance to ensure consistent evaluations of both real and AI-generated images (Figure 6 left). Examples of specific survey questions for cultural contexts, such as those related to Azerbaijan, are shown in Figure 6 right top and right bottom. The estimated time to complete the survey was approximately 1.5 to 2 hours. Table 3 presents the Fleiss’ Kappa scores, which reflect the level of agreement between annotators for each country.

### v2

@@ Abstract
Text-to-image diffusion models have recently enabled the creation of visually compelling, detailed images from textual prompts. However, their ability to accurately represent various cultural nuances remains an open question. In our work, we introduce CultDiff benchmark, evaluating whether state-of-the-art diffusion models can generate culturally specific images spanning ten countries. We show that these models often fail to generate cultural artifacts in architecture, clothing, and food, especially for underrepresented country regions, by conducting a fine-grained analysis of different similarity aspects, revealing significant disparities in cultural relevance, description fidelity, and realism compared to real-world reference images. With the collected human evaluations, we develop a neural-based image-image similarity metric, namely, CultDiff-S, to predict human judgment on real and generated images with cultural artifacts. Our work highlights the need for more inclusive generative AI systems and equitable dataset representation over a wide range of cultures.
@@
Zahra Bayramli1,3thanks:  Equal contribution. Co-first authors:    {zahrabayramly, ayhansuleymanzade}@kaist.ac.krthanks:  Work done during undergraduate studies at the School of Computing, KAIST.   Ayhan Suleymanzade [^1] footnotemark: 1   Na Min An2   Huzama Ahmad2 Eunsu Kim1   Junyeong Park1   James Thorne2thanks:  Corresponding author   Alice Oh133footnotemark: 3 1School of Computing, KAIST  2Graduate School of AI, KAIST  3SOCAR
@@ 1 Introduction
Text-to-image (T2I) diffusion models recently have shown a significant advance in generating high-quality images from text prompts(Saharia et al., 2022; Ramesh et al., 2022; Betker et al., 2023; Podell et al., 2023; Blattmann et al., 2023; Esser et al., 2024). Remarkable success has been demonstrated in producing realistic and detailed visual representations, ranging from generating complex scenes Liu et al. (2024c) to performing style transfer using diffusion models Wang et al. (2023); Zhang et al. (2023). While it is crucial to address the challenge of generating culture-aware imagery, current diffusion models have yet to be evaluated on their ability to create a distribution of accurate and fair representations of images Cao et al. (2024b); Liu et al. (2023).
A recent study, Basu et al. (2023) reveals that while current diffusion models can generate realistic and contextually relevant images, they frequently fail to achieve geographical inclusiveness, generating globally representative images. Furthermore, a critical gap remains in these models’ ability to produce culturally accurate representations, failing to create concepts or artifacts for particular cultures, thus, lacking in culture awareness Kannen et al. (2024). While benchmark datasets have been developed to capture cultural diversity Jha et al. (2023); Myung et al. (2024); Li et al. (2024c), to the best of our knowledge, there is limited prior work exploring the ability of diffusion models to generate culturally specific artifacts, particularly those representing underrepresented cultures Kannen et al. (2024); Nayak et al. (2024).
@@ 2.1 Cultural-Aware Datasets
Beginning with the benchmark to evaluate image quality from simpler text prompts using text-to-image (T2I) diffusion models, such as HPDv2 Wu et al. (2023) using human preferences, more recently advanced efforts include creating comprehensive benchmarks such as T2I-CompBench Huang et al. (2023) and GenAI-Bench Li et al. (2024a). These prior benchmarks assess the performance of T2I models across various aspects, including realism, fidelity, and compositional text-to-image generation. Although a very recent study Jha et al. (2024) has started investigating global representation in T2I models, their emphasis has primarily been on regional social stereotypes. From another perspective, Khanuja et al. (2024) proposed the task of image transcreation for cultural relevance demonstrating that image-editing models struggle to adapt images meaningfully across cultures.
@@
On the other hand, many widely-used datasets for training T2I synthesis models, such as LAION-400M Schuhmann et al. (2021), tend to exhibit Anglo-centric and Euro-centric biases, as noted by Birhane et al. (2021). These biases skew the representation of cultures in generated images, often favoring Western perspectives Mihalcea et al. (2024). In response to this, several researchers have worked to create datasets that better represent diverse cultures. For example, the MaRVL dataset Liu et al. (2021) was specifically designed to include a broader array of languages and cultural concepts, covering regions such as Indonesia, Swahili-speaking Africa, Tamil-speaking South Asia, Turkey, and Mandarin-speaking China and addressing biases in datasets that predominantly focused on North American and Western European cultures. Likewise, Culturally-aware Image Captioning (CIC) framework Yun and Kim (2024) was proposed to enhance cultural diversity in image descriptions.
@@
In the Large Language Models (LLMs) domains, SeeGULL dataset Jha et al. (2023) broadens stereotype benchmarking to encompass global and regional diversity, and BLEnD benchmark Myung et al. (2024) evaluates the cultural knowledge of LLMs across various languages, including low-resource ones. Similarly, the Dollar Street dataset Gaviria Rojas et al. (2022) sought to capture everyday life across a wide variety of socioeconomic backgrounds, presenting a more globally inclusive view. Furthermore, CulText2I Ventura et al. (2023) was introduced as a multilingual benchmark and evaluation framework to analyze cultural encoding in TTI models across languages, cultural concepts, and domains, highlighting how language impacts cultural representation in generated images. In addition, Liu et al. (2023) introduced the CCUB dataset, which was developed to promote cultural inclusivity by collecting images and captions representing the cultural contexts of different countries. Liu et al. (2024a) introduced a Challenging Cross-Cultural (C3) benchmark to evaluate failures in culturally specific generation and proposed fine-tuning on curated data. In the multimodal large language model (MLLM) domain, recent benchmarks such as CVQA Romero et al. (2024) focused on culturally aware question answering by capturing nuances across a diverse set of languages. Most recently, CUBE Kannen et al. (2024) is a large-scale dataset of cultural artifacts spanning 8 countries across different geo-cultural regions for evaluating cultural diversity. Our work further contributes to dataset creation by expanding the focus to include low-resource cultures, thereby addressing gaps between overrepresented and underrepresented cultural contexts.
@@ 3.1 CultDiff Dataset Construction
To acquire real images for training and evaluation, we collected images for each artifact across all categories and countries from the internet using the generated prompts. Recognizing that real-world images vary significantly due to factors such as location, camera angle, lighting, and weather conditions, we aimed to minimize potential biases by selecting five images instead of just one for each artifact. The images were scraped using Bulk Bing Image Downloader (BBID) [^4] https://github.com/ostrolucky/Bulk-Bing-Image-downloader, an open-source Python tool that enables fast downloads and filters out adult content.
@@
3.2 Human Evaluation
A critical part of our experimental design involved human annotation to assess the cultural and contextual relevance of AI-generated images. We first created a survey comprising 150 images for all ten countries (e.g., United States, China, Azerbaijan, and Indonesia), where the images for each country were equally divided across three categories: architecture, clothing, and food items. Each survey included a set of four images for each artifact: three real images (collected via web scraping) and one AI-generated image (produced by one of the three diffusion models).
To prevent model bias, each human annotator was distributed with randomly selected images generated from one of three diffusion models for each question. Hence, the degree of annotator agreement implies varying image quality generated by different models (Appendix A.3 for details). Three human annotators from each country participated in the study, ensuring cultural familiarity and context-specific insights. All 30 annotators were asked to evaluate each image based on the following multiple-choice questions:
Image-image similarity scores
are collected for each category as follows:
Image-text alignment scores
Realism scores
are assessed based on how realistic the AI-generated image appears.
3.3 Model Evaluation
3.3.1 Pair and Dataset Creation
@@ 3.3.2 Similarity Metric
To build a model-based image-image similarity metric, CultDiff-S, we adopted a contrastive learning approach, utilizing Vision Transformers (ViT) (Alexey, 2020) to learn embeddings for both real and AI-generated images, as illustrated in Figure 2. The global attention mechanism in ViT enables effective "local-to-global" correspondence, enhancing self-supervised learning (Caron et al., 2021; Mo et al., 2023). This ViT property is particularly important in our work, where it helps evaluate the cultural and contextual relevance of images by capturing both local features and their broader context. To achieve this, we employed a weighted margin loss during the training process, utilizing the scores (collected from our human annotation process explained in Section 3.2) as weights ($w_{i}$, $i=1,2,...,N$ in Eq. 3.3.2), motivated by previous studies that utilized GPT feedback scores as the weights in building sentence embedding models Cheng et al. (2023); An et al. (2024). These scores are normalized to 0 to 1 and applied to their corresponding real-synthetic (human-annotated) positive and negative image pairs. For real-image pairs that were not annotated by humans, we assigned a default weight of $w=1$. Our margin loss is defined as:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}w_{i}\cdot\Big{(}y_{i}\cdot d_{i}^{2}+(1-y_{i})\cdot\max(0,m-d_{i})^{2}\Big{)},$
@@ 4.1.2 Image-Description Match
Our analysis focused on the responses to the second question (Q2), which assessed the fidelity of the generated content to the provided description. Figure 3 presents the results for the FLUX model’s outputs across architecture, clothing, and food, respectively. The United States and the United Kingdom consistently achieved the highest average fidelity scores across all three categories. In contrast, Azerbaijan, South Korea, and Ethiopia recorded the lowest average scores of 2.28, 1.55, and 1.54 in the architecture, food, and clothing categories, respectively. It is important to note that there is no significant statistical difference between the overrepresented and underrepresented countries in our study, likely due to the small dataset. However, the country rankings remain relevant. In fact, these average scores reveal a noticeable gap between the scores of the United States and those of other countries across all categories. These results suggest a potential bias in the model, favoring countries with a more extensive cultural presence on the internet. Addressing this limitation by incorporating more diverse and balanced datasets could improve the model’s fidelity for underrepresented countries and enhance its overall cultural accuracy Ananthram et al. (2024); Liu et al. (2025).
Results for Stable Diffusion XL and Stable Diffusion 3 Medium Diffusers (see Appendix A.4 for corresponding graphs) also indicate that the United States consistently receives higher scores. Conversely, countries like Azerbaijan and Ethiopia frequently appear in the lower half of the rankings. However, the relative ordering of countries varies across models.
@@ 4.1.3 Realism
4.2 Similarity Metric Analysis
After training the similarity model, we evaluated its performance using our test dataset pairs. For each evaluation pair, we computed additional similarity metrics, including Fréchet Inception Distance (FID) (Heusel et al., 2017), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), and the Structural Similarity Index Measure (SSIM) (Wang et al., 2004). Additionally, we calculated the correlation between image-image similarity scores—derived by averaging the responses to the first four questions in the survey—and each of these similarity metrics. In our metric model, we compute the cosine similarity of the embeddings produced by the model during evaluation as the final similarity score. The results of these correlations are summarized in Table 2. The proposed metric, CultDiff-S, achieves notably higher correlation values, suggesting the usefulness of using our automatic pre-trained image-image metric, showing potential to replace the previous metrics. However, we leave future work to further enhance model correlations with human judgments.
@@ 5.1 Human-Aligned Image-Image Similarity Evaluation Metric
Designing an image-image similarity metric that closely correlates with human rankings is inherently difficult due to the complexity of cultural attributes and subjective interpretations. Culture itself is a broad and multifaceted concept, making it challenging to define in a universally accepted manner Liu et al. (2024b). In fact, many previous studies in related domains that incorporate cultural aspects either provide only a high-level definition or avoid in-depth discussions on its nuances Adilazuarda et al. (2024); Huang and Yang (2023); Zhou et al. (2023); Cao et al. (2024a); Wan et al. (2023). In our study, we treated countries as representatives of distinct cultures. However, individuals, even within the same country, may have diverse interpretations of cultural artifacts, reflecting regional, ethnic, and personal differences in cultural perception, which can lead to subjective similarity scores given by human annotators. Additionally, cultural artifacts often evolve over time, influenced by various internal or external factors, further complicating the task of cultural representations. As a result, an effective similarity metric that correlates more closely with human perception must account for both objective attributes, such as structural and visual features, and subjective factors, including historical context, regional variations, and personal interpretations. We believe that developing such a metric necessitates integrating human feedback into the process, combining both data-driven approaches and human-centered evaluations to bridge the gap, as has been demonstrated in other contexts and domains in prior works Xu et al. (2023); Stiennon et al. (2020); Ouyang et al. (2022), as well as in our research.
@@ 6 Conclusion
Acknowledgments
This research was supported by the MSIT (Ministry of Science, ICT), Korea, under the National Program for Excellence in SW), supervised by the IITP (Institute of Information & communications Technology Planing & Evaluation) in 2024 (2022-0-01092) and National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.RS-2024-00406715), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509258 and No. RS-2024-00469482, Global AI Frontier Lab).
This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program through which leading foundation models hosted by Microsoft Azure along with access to Azure credits were provided to conduct the research.
@@ A.1 Prompt Details
1. “A panoramic view of {landmark} in {country}, realistic”
2. “An image of {clothes} from {country} clothing, realistic”
3. “An image of {food} from {country} cuisine, realistic”
@@
1. “A panoramic view of the Empire State Building in the United States, realistic”
2. “An image of Hanfu from Chinese clothing, realistic”
3. “An image of plov from Azerbaijani cuisine, realistic”
@@
A.2 Model Parameters
A.2.1 Image Generation
The models used for image generation have the following parameter counts: Stable Diffusion XL Podell et al. (2023) has 2.6 billion U-Net parameters, Stable Diffusion 3 Medium Esser et al. (2024) has 2 billion parameters, and FLUX.1-dev Labs (2023) is a 12-billion-parameter rectified flow transformer.
A.2.2 Model Training
For model training, we used ViT-Base Alexey (2020), which has 86 million parameters, 12 layers, a hidden size of 768, an MLP size of 3072, and 12 attention heads. We trained our contrastive learning model using ViT-Base for 10 epochs with a learning rate of $1\times 10^{-4}$, a batch size of 32, and an image resolution of $224\times 224$.
A.3 Survey Question Details
This section provides additional details regarding the survey questions and instructions used during the study. Annotators received clear guidance to ensure consistent evaluations of both real and AI-generated images (Figure 6 left). Examples of specific survey questions for cultural contexts, such as those related to Azerbaijan, are shown in Figure 6 right top and right bottom. The estimated time to complete the survey was approximately 1.5 to 2 hours. Table 3 presents the Fleiss’ Kappa scores, reflecting annotator agreement across different diffusion models (Stable Diffusion XL, Stable Diffusion 3 Medium Diffusers, FLUX). The relatively low agreement scores may stem from the fact that annotators from the same country evaluated images generated by different models rather than assessing a shared set of images. Note that we showed a comparative analysis across three models in terms of image-description match (Section 4.1.2) and realism (Section 4.1.3).